A metal has Er= 9, Wp=7×10^15 rad s^-115, v=2×10^12 s^-1 . Calculate the skin depth for microwave of 4 cm vacuum wavelength.
1
Expert's answer
2019-04-17T11:41:17-0400

The frequency is given by


"f=\frac {c}{\lambda} (1)"

where λ=4 cm

Using (1) we got: f=7.5GHz


Skin depth δ is given by formula


"\delta= \sqrt{\frac {\rho}{\pi f \mu}} (2)"

where ρ is the resistivity of material, f is the frequency, μ is the permeability

Using (2) we got:

Skin Depth for Copper is equal to 0.7528 μm